Show the COMPLETE solution on the given problems.

1. Evaluate the line integral ∮2
1
Expert's answer
2020-03-26T13:04:06-0400

Evaluate the line integral "\oint_C2ds" where "C" is the unit circle.


"\oint_Cf(x, y)ds=\displaystyle\int_{a}^bf(x(t),y(t))\sqrt{\big({dx \over dt}\big)^2+\big({dy \over dt}\big)^2}dt"

We first need parametric equations to represent "C." The unit circle can be parametrized by means of the equations


"x=\cos t, y=\sin t, \ 0\leq t\leq2\pi."

"{dx \over dt}=-\sin t,\ {dy \over dt}=\cos t"

"\sqrt{\big({dx \over dt}\big)^2+\big({dy \over dt}\big)^2}=\sqrt{\big(-\sin t\big)^2+\big(\cos t\big)^2}=1"

"\oint_C2ds=\displaystyle\int_{0}^{2\pi}2(1)dt=2\big[t\big]\begin{matrix}
 2\pi \\
 0
\end{matrix}=2(2\pi-0)=4\pi"